Could you help me solve 3/y + 6/5
Answer by checkley71(8403)

3/Y+6/5=0
3/Y=-6/5
-6Y=15
Y=15/-6
Y=-2.5 SOLUTION
PROOF
3/2.5=6/5
6*2.5=3*5
15=15